A painting crew can paint a structure in 12 hr or in 7.2 Jr when working with a second crew. How long would it take the second crew to do the job if working alone


1
Expert's answer
2021-11-03T12:22:15-0400

It fololws that a first painting crew can paint 112\frac{1}{12} part of a structure in 1 hr, and they both can paint 17.2=536\frac{1}{7.2}=\frac{5}{36} part of a structure in 1 hr. Therefore, the second crew can paint 536112=5336=236=118\frac{5}{36}-\frac{1}{12}=\frac{5-3}{36}=\frac{2}{36}=\frac{1}{18} part of a structure in 1 hr. Therefore, the second crew do the job if working alone 18 hours.